Quick Meal is a fast food outlet in Mona that sells box lunches for $120 each. Its total cost function is TC = 80,000 + 20Q + 0.02Q2

a. Should the owners close the restaurant? Explain!

b. What is the profit /loss of the business?
Answer #1

Answer

a)

No.

A firm shuts down if the price is below minimum average varible costs.

the firm produces at MC=P

MC=change in total cost =dTC/dQ=20+0.04Q

equating the MC=P

20+0.04Q=120

0.04Q=100

Q=2500

AVC=VC/Q

VC=20Q+0.02Q^2 ................ the TC except for constant term

AVC=20+0.02Q

AVC=20+0.02*2500

AVC=70

the P>AVC so the firm can produce to minimize losses or to maximize profit.

==============

b)

Profit=TR-TC

TR=P*Q=2500*120=300000

TC=80000+20*2500+0.02*2500^2=255000

Profit=300000-255000

=45000

the economic profit is $45000